OpenStudy (anonymous):

Solve the following equation by transforming it into a perfect square trinomial: 3x^2 + 18x = 21

OpenStudy (nurali):

3x² + 18x = 21 x² + 6x = 7 x² + 3x = 7 + 3² x² + 3x = 7 + 9 (x + 3)² = 16 x + 3 = 16 x = 16 - 3, x = 13 x = - 16 - 3, x = - 19 Answer: (- 19, 13)

OpenStudy (anonymous):

thank you

OpenStudy (nurali):

Anytime.